OpenStudy (anonymous):

0^0 is 0 or the solution doesnt exist. why??

OpenStudy (anonymous):

?

OpenStudy (anonymous):

zero to the power of zero. equals 0 or it doesn't exist?

OpenStudy (akshay_budhkar):

the answer is doesn't exist

OpenStudy (anonymous):

0^0 = 1

OpenStudy (akshay_budhkar):

you can verify from here http://www.wolframalpha.com/input/?i=0^0

OpenStudy (anonymous):

it is equal to 1 anythng to the power of zero is 1

OpenStudy (anonymous):

i need an argument why or why not.. hmmm...

OpenStudy (anonymous):

yes... 0^0 in terms of a determinate is indeterminate

OpenStudy (akshay_budhkar):

anything to the power 0 is 1 other than 0 itself ofcourse.. if you see the graph of x^0 it is discontinuous at x=0

jimthompson5910 (jim_thompson5910):

0 to any power is 0 anything to the power of 0 is 1 see the conflict? So 0^0 is indeterminate

OpenStudy (anonymous):

there is conflict but wouldn't that mean the theory itselt is a conflict lol?

jimthompson5910 (jim_thompson5910):

more formally, \[\Large 0^x=0\] where x is nonzero \[\Large x^0=1\] where x is nonzero

OpenStudy (anonymous):

0=1 is a false staemtn

jimthompson5910 (jim_thompson5910):

exactly, so saying 0^0 = 1 is not correct, neither is saying 0^0 = 0

OpenStudy (anonymous):

lol this question is a puzzle itself?

OpenStudy (akshay_budhkar):

Ah so that is what i am telling.. it does not exist or else if we have x tending to zero then we have to evaluate the limit using the most appropriate method availbale

jimthompson5910 (jim_thompson5910):

i guess you can see it that way, but you'll find that x=0 isn't in the domain of x^x

OpenStudy (akshay_budhkar):

it is one of the seven to eight indeterminate forms.. it is a part of limits

OpenStudy (akshay_budhkar):

So is your query cleared? or did it get more intensified?

OpenStudy (anonymous):

Def confused, but it's ok, haha

OpenStudy (akshay_budhkar):

Why are you confused??

OpenStudy (anonymous):

what kind of math are you taking? like is this a college level course? because this kind of question is a little bit of a joke and yet a serious question in my ibh maths class and well, i wouldn't say there's no answer, i'd just say that it's undefined. if you graph an xˆy function, its a function of 2 variables and so it's not continuous in origin. but in the case of 0ˆ0, thinks get really messy.. i remember i asked my math teacher this question once and all he did was he looked at me and said, "because mathematicians said so."

OpenStudy (anonymous):

it's a college level math course but the teacher is obviously trying to see what i'm going to answer. we have to explain why we think it's 0 or why we think there is no solution